OpenStudy (dj3strella):

Help 4. Divide. (3a3 + 6a2) ÷ 3a (Points : 4) 3a^3 + 6a^2 – 3a 3a^3 + 6a^2 + 3a a^2 + 2a a^2 + 6a^2

ganeshie8 (ganeshie8):

\(\large \frac{3a^3 + 6a^2}{3a}\)

ganeshie8 (ganeshie8):

\(\large \frac{3a^3 }{3a} + \frac{6a^2}{3a}\)

ganeshie8 (ganeshie8):

3a cancels out in first term, and in second term also, eh ?

ganeshie8 (ganeshie8):

\(\large \frac{3a \times a^2 }{3a} + \frac{3a \times 2a}{3a}\) \(\large a^2 + 2a\)

OpenStudy (dj3strella):

thank you